# The Race between Technology and Woman:

### v8

@@ 3.1.3 Equilibrium
We consider the competitive equilibrium that satisfies conditions (ii) and (iii) throughout this paper. When we extend this model to allow for endogenous labor supply and imperfect labor mobility, we modify condition (i) or (iv) as described in Section 3.5.
@@ 3.2 Aggregate elasticity of substitution
where $\ell_{f}=\sum_{n=1}^{N}\ell_{fn}$ for $f\in\{1,\ldots,F\}$. [^3] The Morishima elasticity of substitution is a measure of curvature of the isoquant in line with the original concept by Hicks (1932) and a sufficient statistic to measure the contributions of factor prices to relative factor shares, as noted by Blackorby and Russell (1989). The partial derivative of the log of the production (cost) function with respect to the log of the factor quantity (price) equals the factor share of income in the aggregate economy, defined as $\Lambda_{\ell_{f}}=w_{f}\ell_{f}/py$. The aggregate elasticities can be measured by calculating the extent to which the factor shares of income vary according to the quantity or price of factors. Throughout this and next two subsections, we suppress the time subscript $t$ for notational simplicity.
Proposition 1.
@@
Appendix A.1 provides the proofs of all propositions.
@@
The aggregate elasticity of substitution (14) can be decomposed into two parts:
$\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=\left[\sum_{n=1}^{N}\left(\frac{\ell_{fn}}{\ell_{f}}\right)\frac{\partial\ln\lambda_{\ell_{f}n}}{\partial\ln w_{g}}-\sum_{n=1}^{N}\left(\frac{\ell_{gn}}{\ell_{g}}\right)\frac{\partial\ln\lambda_{\ell_{g}n}}{\partial\ln w_{g}}+1\right]+\left[\sum_{n=1}^{N}\left(\frac{\ell_{fn}}{\ell_{f}}-\frac{\ell_{gn}}{\ell_{g}}\right)\frac{\partial\ln\zeta_{n}}{\partial\ln w_{g}}\right],$ (15)
where $\partial\ln\zeta_{n}/\partial\ln w_{g}=\partial\ln(c_{n}/c)/\partial\ln w_{g}+\partial\ln(p_{n}/p)/\partial\ln w_{g}=[1-1/(1-\eta)](\lambda_{\ell_{g}n}-\Lambda_{\ell_{g}})$.
The first square bracket captures the extent to which the producer’s optimal mix of factors changes in response to relative factor prices. The elasticity of substitution in the sectoral cost function is defined as $\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}=\partial\ln[(\partial\mathcal{C}_{n}/\partial w_{f})/(\partial\mathcal{C}_{n}/\partial w_{g})]/\partial\ln(w_{g}/w_{f})$. If factor intensities are equal across sectors (i.e., $\ell_{fn}/\ell_{gn}=\ell_{fn^{\prime}}/\ell_{gn^{\prime}}=\ell_{f}/\ell_{g}$ for $n\neq n^{\prime}$), the aggregate elasticity (15) is the weighted average of the sectoral elasticities; that is, $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=\sum_{n=1}^{N}(\ell_{fn}/\ell_{f})\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}$, where the weight is the quantity share of factor $f$ in each sector. Furthermore, if the elasticities of substitution are equal across sectors (i.e., $\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}=\epsilon_{\ell_{f}\ell_{g}n^{\prime}}^{\mathcal{C}}$ for $n\neq n^{\prime}$), the aggregate elasticity coincides with the sectoral elasticity. The second square bracket captures the extent to which the consumer’s optimal mix of goods (i.e., the size of each sector) changes in response to relative factor prices. An increase in the factor price lowers the consumption of goods, while it raises the price of goods due to an increase in the marginal cost of production. When the demand is inelastic (i.e., 1/($11/(1-\eta)<$ 1), the effect of the increased price on the expenditure share exceeds that of the decreased consumption. In that case, an increase in the price of factor $g$ raises the size of sectors that are intensive in factor $g$ more than the average sector (i.e., $\lambda_{\ell_{g}n}>\Lambda_{\ell_{g}}$). Consequently, an increase in the price of factor $g$ raises the equilibrium quantity of factor $f$ relative to factor $g$ in sectors more intensive in factor $f$ than in factor $g$ (i.e., $\ell_{fn}/\ell_{f}>\ell_{gn}/\ell_{g}$). The aggregate elasticity can be expressed in a simpler form in the case of two inputs without a functional form assumption on the sectoral production function.
Corollary 2.
@@ 3.3.1 Labor shares
The aggregate elasticities of substitution, $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}$ ($\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$), are sufficient statistics to measure the contributions of factor quantities (prices) to the factor shares of income if the factor shares of income are known. We denote by $\Delta\ln z$ the log change of variable $z$ from one period to another.
@@
We refer to the first two terms in equation (16) as the relative quantity effect, the first two terms in equation (17) as the relative price effect, and the third term in equations (16) and (17) as the factor complementarity effect. The direction of the relative price and quantity effects on the income share of factor $f$ depends on whether the elasticity of substitution between factors $f$ and $g$ exceeds one, whereas the direction of the factor complementarity effect on the income share of factor $f$ depends on whether factor $g$ is more complementary to factor $f$ than to factor $h$. If only one type of capital and one type of labor are used to produce goods and services (or equivalently, if all types of labor are perfectly substitutable, and all types of capital are perfectly substitutable), the factor complementarity effect is absent. In such a case, the direction and magnitude of changes in the labor share are governed by the elasticity of substitution between capital and labor and the degree of capital deepening, as first noted by Hicks (1932).
Corollary 3.
or
As is well known, in this simple case, the labor share can decline with a rise in the capital–labor ratio or the wage–rental price ratio if $\epsilon>1$. Karabarbounis and Neiman (2014) attribute the decline in the labor share to capital-embodied technological change on the basis of findings that the estimated elasticity of substitution exceeds one and that the rental price of capital fell in many countries. Elsby, Hobijn and Şahin (2013) attribute it to other factors on the basis of the observation that the timing of decline in the labor share did not necessarily align with the timing of capital deepening in the United States.
@@
where $\Delta\ln\Lambda_{\ell_{f}}$ is given by equation (16) or (17), and the weight is the wage bill share of each type of labor. Given the fact that the majority of labor is male unskilled, the largest weight is assigned to the rate of change in the income share of male unskilled labor. Whether the labor share can decline with progress in ICT depends predominantly on the extent to which the income share of male unskilled labor can decline with a fall (rise) in the rental price (quantity) of ICT capital. If the aggregate elasticity of substitution between ICT capital and male unskilled labor exceeds one, a fall (rise) in the rental price (quantity) of ICT capital can serve to reduce the labor share. Moreover, if ICT capital is more substitutable with male unskilled labor than with other types of labor, a fall (rise) in the price (quantity) of ICT capital can further reduce the labor share.
@@ 3.3.2 Relative wages
The aggregate elasticities of substitution, $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}$ ($\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$), are sufficient statistics to measure the contributions of factor quantities (prices) to the relative wages of (relative demand for) different types of labor.
@@
We refer to the first term as the factor complementarity effect and the second and third terms in equation (19) as the relative quantity effect again. The direction of the factor complementarity effect on the wages of factor $f$ relative to factor $g$ depends on whether factor $h$ is more complementary to factor $f$ than to factor $g$, whereas the direction of the relative quantity effect depends on whether each factor increases. Equation (19) is a multi-sector, multi-factor general equilibrium version of the model of the race between technology and skills, which nests the standard models of skill differentials.
Corollary 4.
Corollary 5.
Suppose that production requires two types of capital (capital equipment, $k_{e}$, and capital structures, $k_{s}$) and two types of labor (skilled labor, $\ell_{h}$, and unskilled labor, $\ell_{u}$), and the aggregate elasticities of substitution satisfy $\epsilon_{\ell_{h}k_{s}}^{\mathcal{F}}=\epsilon_{\ell_{u}k_{s}}^{\mathcal{F}}=1$ and $1/\epsilon_{\ell_{u}\ell_{h}}^{\mathcal{F}}-1/\epsilon_{\ell_{h}\ell_{u}}^{\mathcal{F}}=1/\epsilon_{\ell_{h}k_{e}}^{\mathcal{F}}-1/\epsilon_{\ell_{u}k_{e}}^{\mathcal{F}}$. Equation (19) becomes the Krusell, Ohanian, Rios-Rull and Violante (2000) model:
This equation implies that the skill wage gap can increase with a rise in capital equipment if capital equipment is more complementary to skilled labor than to unskilled labor (i.e., $\epsilon_{\ell_{u}k_{e}}^{\mathcal{F}}>\epsilon_{\ell_{h}k_{e}}^{\mathcal{F}}$). This effect is sufficiently large to account for a rise in the skill premium in OECD countries when ICT capital is distinguished from non-ICT capital (Taniguchi and Yamada, 2022).
@@
Equation (19) shows how gender and skill premia can vary according to the shift in demand due to progress in ICT and the shift in supply due to advances in educational attainment and female employment. The first term captures the shift in demand. The wages of factor $f$ relative to factor $g$ can increase with a rise in factor $h$ if factor $h$ is more complementary to factor $f$ than to factor $g$ (i.e., $\epsilon_{\ell_{g}\ell_{h}}^{\mathcal{F}}>\epsilon_{\ell_{f}\ell_{h}}^{\mathcal{F}}$). In particular, if ICT capital is more complementary to skilled (female) labor than to unskilled (male) labor, the skill (gender) wage gap increases (decreases) with a rise in ICT capital. The magnitude of this factor complementarity effect is directly proportional to the difference between $\epsilon_{\ell_{g}k_{i}}^{\mathcal{F}}$ and $\epsilon_{\ell_{f}k_{i}}^{\mathcal{F}}$, as well as the rate of increase in $k_{i}$. The second and third terms capture the shift in supply. The wages of factor $f$ relative to factor $g$ can decrease with a rise (fall) in the quantity of factor $f$ ($g$). The magnitude of this relative quantity effect is inversely proportional to $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}$ and $\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{F}}$ and directly proportional to the rate of increase in $\ell_{f}$ relative to $\ell_{g}$. Given the fact that the rate of increase in skilled (female) labor is greater than that in unskilled (male) labor, the relative quantity effect should be negative (positive) for the skilled–unskilled (male–female) wage gap. Consequently, the factor complementarity effect and the relative quantity effect would work in opposite directions. Whether the skill (gender) wage gap eventually widens (narrows) depends on the relative magnitude of these two effects.
@@ 3.4 Effects of technological change
The analysis thus far is useful for decomposing the sources of changes in the relative wages and income shares of different types of labor. We can measure the extent to which each factor contributes to changes in relative wages and labor shares using equations (16) and (19) for given values of the aggregate elasticities of substitution. However, if the equilibrium prices and quantities of factors can change in response to technological change, those equations are insufficient for determining the effects of technological change. Here, we derive the analytical expressions for changes in the equilibrium relative wages and income shares of different types of labor due to capital-embodied technological change and sector-specific disembodied technological change. Throughout this and next subsections, we focus on steady-state equilibria in which all variables are constant over time and examine changes before and after information revolution. We denote by $d\ln z$ the log change of variable $z$ from one steady state to another.
@@
where $\Upsilon_{\left(f,g\right)}^{q}$ is the $fg$-th element of the $F_{\ell}\times F_{k}$ matrix $\Upsilon^{q}$, and $\Upsilon_{\left(f,n\right)}^{A}$ is the $fn$-th element of the $F_{\ell}\times N$ matrix $\Upsilon^{A}$. The matrices $\Upsilon^{q}$ and $\Upsilon^{A}$ are described in Appendix A.1. Changes in labor shares due to technological change are given by
@@
Equations (21) and (22) express how the relative wages and income shares of different types of labor can change in response to capital-embodied technological change and sector-specific disembodied technological change in a multi-sector, multi-factor general equilibrium model of investment-specific technological change. The effect of capital-embodied technological change on relative wages, which is represented by the first term in equation (21), is related to the factor complementarity effect in equation (19). The effect of capital-embodied technological change on labor shares, which is represented by the first term in equation (22), is related to the relative price effect and the factor complementarity effect in equation (17). The first term in equation (22) can be divided into two parts, one in the first line and the other in the second line. The two parts capture the effects through changes in wages and rental prices, respectively. Most elements of the matrix $\Upsilon^{q}$ consist of the aggregate elasticities of substitution $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$, and the others consist of the factor shares of income $\Lambda_{\ell_{f}}$, as detailed in the appendix. Equations (21) and (22) show that the aggregate elasticities of substitution are sufficient statistics to quantify the effects of technological change on the relative wages and income shares of different types of labor if the income shares of each factor are known.

### v9

@@ 3.1.3 Equilibrium
We consider the competitive equilibrium that satisfies conditions (ii) and (iii) throughout this paper. When we extend this model to allow for endogenous labor supply or imperfect labor mobility, we modify condition (i) or (iv) as described in Appendix A.
@@ 3.2 Aggregate elasticity of substitution
where $\ell_{f}=\sum_{n=1}^{N}\ell_{fn}$ for $f\in\{1,\ldots,F\}$. [^4] The Morishima elasticity of substitution is a measure of curvature of the isoquant in line with the original concept by Hicks (1932) and a sufficient statistic for measuring the contributions of factor prices to relative factor shares, as noted by Blackorby and Russell (1989). The partial derivative of the log of the production (cost) function with respect to the log of the factor quantity (price) equals the factor share of income in the aggregate economy, defined as $\Lambda_{\ell_{f}}=w_{f}\ell_{f}/py$. The aggregate elasticities of substitution can be derived by looking at the extent to which the factor shares of income vary according to the quantity or price of factors. Throughout this and next two subsections, we suppress the time subscript $t$ for notational simplicity. The following lemma recasts the results of Baqaee and Farhi (2019) for value-added output. [^5] Baqaee and Farhi (2019) derive expressions for the elasticities of substitution in the aggregate production and cost functions when a micro-level production function takes a nested CES form, and firms are connected through input–output linkages (Propositions 3 and 8). In the appendix, they refine these expressions without imposing any functional form on the micro-level production function. Oberfield and Raval (2021) derive expressions for the elasticity of substitution in the aggregate cost function both when the micro-level production function takes a nested CES form (Proposition 2) and when no functional form is imposed on the micro-level production function (Proposition 2’).
Lemma 1.
@@
Appendix B contains the proofs of this lemma and all propositions. The lemma and propositions apply not only to sector-level production functions but also to firm- and establishment-level production functions.
@@
The aggregate elasticity of substitution (13) can be expressed as a weighted average of the elasticities of substitution in production and consumption:
$\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=\left[\sum_{n=1}^{N}\left(\frac{\ell_{fn}}{\ell_{f}}\right)\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}+\sum_{n=1}^{N}\left(\frac{\ell_{gn}}{\ell_{g}}-\frac{\ell_{fn}}{\ell_{f}}\right)\sum_{f\neq g}^{F}\lambda_{\ell_{f}n}\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}\right]+\left[\sum_{n=1}^{N}\left(\frac{\ell_{gn}}{\ell_{g}}-\frac{\ell_{fn}}{\ell_{f}}\right)\lambda_{\ell_{g}n}\frac{1}{1-\eta}\right],$ (14)
where the elasticity of substitution in the sectoral cost function is defined as $\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}=\partial\ln[(\partial\mathcal{C}_{n}/\partial w_{f})$ $(\partial\mathcal{C}_{n}/\partial w_{g})]/\partial\ln(w_{g}/w_{f})$.
The first square bracket captures the extent to which the producer’s optimal mix of factors changes in response to relative factor prices. If factor intensities are equal across sectors (i.e., $\ell_{fn}/\ell_{gn}=\ell_{fn^{\prime}}/\ell_{gn^{\prime}}=\ell_{f}/\ell_{g}$ for $n\neq n^{\prime}$), the aggregate elasticity (14) is the weighted average of the sectoral elasticities; that is, $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}=\sum_{n=1}^{N}(\ell_{fn}/\ell_{f})\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}$, where the weight is the quantity share of factor $f$ in each sector. Furthermore, if the elasticities of substitution are equal across sectors (i.e., $\epsilon_{\ell_{f}\ell_{g}n}^{\mathcal{C}}=\epsilon_{\ell_{f}\ell_{g}n^{\prime}}^{\mathcal{C}}$ for $n\neq n^{\prime}$), the aggregate elasticity coincides with the sectoral elasticity. The second square bracket captures the extent to which the consumer’s optimal mix of goods (i.e., the size of each sector) changes in response to relative factor prices. An increase in the factor price raises the price of goods due to an increase in the marginal cost of production, which in turn reduces consumption. An increase in the price of factor $g$ reduces the equilibrium quantity of labor $g$ relative to factor $f$ (or equivalently raises the equilibrium quantity of factor $f$ relative to factor $g$) in sectors more intensive in factor $g$ than in factor $f$ (i.e., $\ell_{fn}/\ell_{f}<\ell_{gn}/\ell_{g}$). [^6] Our representation in equation (14) differs from Baqaee and Farhi (2019), who consider a production function with capital, labor, and intermediate inputs and underscore input–output linkages, and from Oberfield and Raval (2021), who consider a production function with capital, labor, and material inputs and focus on the elasticity of substitution between capital and labor.
The aggregate elasticities (12) and (13) can be expressed in a simple form if production requires only one type of capital and one type of labor, as described in Corollary 5 in Appendix C.
@@ 3.3.1 Labor shares
The aggregate elasticities of substitution, $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}$ ($\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$), are sufficient statistics for measuring the contributions of factor quantities (prices) to the factor shares of income if the factor shares of income are known. We denote by $\Delta\ln z$ the log change of variable $z$ from one period to another.
@@
We refer to the first two terms in equation (15) as the relative quantity effect, the first two terms in equation (16) as the relative price effect, and the third term in equations (15) and (16) as the factor complementarity effect. The direction of the relative price and quantity effects on the income share of factor $f$ depends on whether the elasticity of substitution between factors $f$ and $g$ exceeds one, whereas the direction of the factor complementarity effect on the income share of factor $f$ depends on whether factor $g$ is more complementary to factor $f$ than to factor $h$.
If production requires only one type of capital and one type of labor (or equivalently, if all types of labor are perfectly substitutable, and all types of capital are perfectly substitutable), the factor complementarity effect may disappear. In this case, the direction and magnitude of changes in the labor share are determined by the elasticity of substitution between capital and labor and the degree of capital deepening, as described in Corollary 6 in Appendix C.
@@
where $\Delta\ln\Lambda_{\ell_{f}}$ is given by equation (15), and the weight is the wage bill share of each type of labor. Given the fact that the majority of labor is male unskilled, the largest weight is assigned to the rate of change in the income share of male unskilled labor.
@@ 3.3.2 Relative wages
The aggregate elasticities of substitution, $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}$ ($\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$), are sufficient statistics for measuring the contributions of factor quantities (prices) to the relative wages of (relative demand for) different types of labor.
@@
We refer to the first term as the factor complementarity effect and the second and third terms in equation (18) as the relative quantity effect again. The direction of the factor complementarity effect on the wages of factor $f$ relative to factor $g$ depends on whether factor $h$ is more complementary to factor $f$ than to factor $g$, whereas the direction of the relative quantity effect depends on whether each factor increases. Equation (18) is a multi-sector, multi-factor general equilibrium version of the model of the race between technology and skills, which nests the standard models of skill differentials, as described in Corollaries 7 and 8 in Appendix C.
@@
Equation (18) shows how gender and skill premia can vary according to the shift in demand due to progress in ICCT and the shift in supply due to advances in educational attainment and female employment. The first term captures the shift in demand. The wages of factor $f$ relative to factor $g$ can increase with a rise in factor $h$ if factor $h$ is more complementary to factor $f$ than to factor $g$ (i.e., $\epsilon_{\ell_{g}\ell_{h}}^{\mathcal{F}}>\epsilon_{\ell_{f}\ell_{h}}^{\mathcal{F}}$). In particular, if ICCT capital is more complementary to skilled (female) labor than to unskilled (male) labor, the skill (gender) wage gap increases (decreases) with a rise in ICCT capital. The magnitude of this factor complementarity effect is directly proportional to the difference between $\epsilon_{\ell_{g}k_{i}}^{\mathcal{F}}$ and $\epsilon_{\ell_{f}k_{i}}^{\mathcal{F}}$, as well as the rate of increase in $k_{i}$. The second and third terms capture the shift in supply. The wages of factor $f$ relative to factor $g$ can decrease with a rise (fall) in the quantity of factor $f$ ($g$). The magnitude of this relative quantity effect is inversely proportional to $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{F}}$ and $\epsilon_{\ell_{g}\ell_{f}}^{\mathcal{F}}$ and directly proportional to the rate of increase in $\ell_{f}$ relative to $\ell_{g}$. Given the fact that the rate of increase in skilled (female) labor is greater than that in unskilled (male) labor, the relative quantity effect should be negative (positive) for the skilled–unskilled (male–female) wage gap. Consequently, the factor complementarity effect and the relative quantity effect would work in opposite directions. Whether the skill (gender) wage gap eventually widens (narrows) depends on the relative magnitude of these two effects.
When production requires four or more types of labor, it may be of interest to examine the sources of change in the average wage for one group relative to that for another group. The average wage of labor $h$ in group $H$ is defined as $\overline{w}_{H}=\sum_{h\in H}(\ell_{h}/\sum_{h\in H}\ell_{h})w_{h}$. The rate of change in the average wage for group $H$ relative to that for group $U$ can then be expressed as
$\Delta\ln\left(\frac{\overline{w}_{H}}{\overline{w}_{U}}\right)=\sum_{h\in H}\left(\frac{w_{h}\ell_{h}}{\sum_{h\in H}w_{h}\ell_{h}}\Delta\ln\Lambda_{\ell_{h}}-\frac{\ell_{h}}{\sum_{h\in H}\ell_{h}}\Delta\ln\ell_{h}\right)\\ -\sum_{u\in U}\left(\frac{w_{u}\ell_{u}}{\sum_{u\in U}w_{u}\ell_{u}}\Delta\ln\Lambda_{\ell_{u}}-\frac{\ell_{u}}{\sum_{u\in U}\ell_{u}}\Delta\ln\ell_{u}\right),$ (20)
where $\Delta\ln\Lambda_{\ell_{f}}$ is given by equation (15). The rate of change in the ratio of average wages between two groups equals the difference between the weighted average of the rates of change in the income shares of different types of labor, with each type weighted by its wage bill share, and the weighted average of the rates of change in the quantities of different types of labor, with each type weighted by its quantity share.
@@ 3.4 Effects of technological change
The analysis thus far is useful for isolating the sources of changes in the relative wages and income shares of different types of labor. We can quantify the extent to which each factor contributes to changes in relative wages and labor shares using equations (15) and (18) for given values of the aggregate elasticities of substitution. However, if the equilibrium prices and quantities of factors can change in response to technological change, those equations are insufficient to fully capture the effects of technological change. Here, we derive the analytical expressions for changes in the equilibrium relative wages and income shares of different types of labor due to capital-embodied technological change and sector-specific disembodied technological change. Throughout this and next subsections, we focus on steady-state equilibria in which all variables are constant over time and examine changes before and after information revolution. We denote by $d\ln z$ the log change of variable $z$ from one steady state to another.
@@
where $\Upsilon_{\left(f,g\right)}^{q}$ is the $fg$-th element of the $F_{\ell}\times F_{k}$ matrix $\Upsilon^{q}$, and $\Upsilon_{\left(f,n\right)}^{A}$ is the $fn$-th element of the $F_{\ell}\times N$ matrix $\Upsilon^{A}$. The matrices $\Upsilon^{q}$ and $\Upsilon^{A}$ are described in Appendix B. Changes in labor shares due to technological change are given by
@@
Equations (21) and (22) express how the relative wages and income shares of different types of labor can change in response to capital-embodied technological change and sector-specific disembodied technological change in a multi-sector, multi-factor general equilibrium model of investment-specific technological change. The effect of capital-embodied technological change on relative wages, which is represented by the first term in equation (21), is related to the factor complementarity effect in equation (18). The effect of capital-embodied technological change on labor shares, which is represented by the first term in equation (22), is related to the relative price effect and the factor complementarity effect in equation (16). The first term in equation (22) can be divided into two parts, one in the first line and the other in the second line. The two parts capture the effects through changes in wages and rental prices, respectively. Most elements of the matrix $\Upsilon^{q}$ consist of the aggregate elasticities of substitution $\epsilon_{\ell_{f}\ell_{g}}^{\mathcal{C}}$, and the others consist of the factor shares of income $\Lambda_{\ell_{f}}$. Equations (21) and (22) show that the aggregate elasticities of substitution are sufficient statistics for quantifying the effects of technological change on the relative wages and income shares of different types of labor if the income shares of each factor are known.